# Reward Centering

### v1

@@ Abstract
We show that discounted methods for solving continuing reinforcement learning problems can perform significantly better if they center their rewards by subtracting out the rewards’ empirical average. The improvement is substantial at commonly used discount factors and increases further as the discount factor approaches one. In addition, we show that if a problem’s rewards are shifted by a constant, then standard methods perform much worse, whereas methods with reward centering are unaffected. Estimating the average reward is straightforward in the on-policy setting; we propose a slightly more sophisticated method for the off-policy setting. Reward centering is a very general idea, so we expect almost every reinforcement-learning algorithm to benefit by the addition of reward centering.
@@
Figure 1: Learning curves showing the difference in performance of Q-learning with and without reward centering for different discount factors on the classic Access-Control Queuing problem. Plotted is the average per-step reward obtained by the agent across 50 runs w.r.t. the number of time steps of interaction. The shaded region denotes one standard error. Details in Section 4.
In this paper, we show that the simple idea of estimating and subtracting the average reward from the observed rewards can lead to a significant improvement in performance when using standard discounting methods such as TD-learning (Sutton, 1988a) or Q-learning (Watkins & Dayan, 1992). The underlying theory dates back to 1962 with Blackwell’s seminal work on dynamic programming in discrete Markov decision processes (MDPs). We are still realizing some of its deeper implications, and we shall discuss two in particular.
@@
    Furthermore, mean-centering the rewards (unsurprisingly) makes standard methods robust to any constant offset in the rewards. This can be useful when applying RL algorithms in problems where the properties of the reward signal are unknown or changing.
We begin with what reward centering is and why it can be beneficial (Section 1). We then show how reward centering can be done, starting with the simplest form (within the prediction problem), and show that it can be highly effective when used with discounted-reward algorithms such TD-learning (Section 2). However, the off-policy setting requires more sophistication, and we propose another way of reward centering based on recent advances in the average-reward formulation for RL (Section 3). Next, we perform a case study of using reward centering with Q-learning, in which we (a) present a convergence result based on recent work by Devraj and Meyn (2021), and (b) showcase consistent trends across a series of control problems that require tabular, linear, and non-linear function approximation (Section 4). Finally, we discuss the limitations of the proposed methods and propose directions of future work (Section 5).
@@ 1 Theory of Reward Centering
We formalize the interaction between the agent and the environment by a finite MDP ${\mathcal{M}}\doteq({\mathcal{S}},{\mathcal{A}},{\mathcal{R}},p)$, where ${\mathcal{S}}$ denotes the set of states, ${\mathcal{A}}$ denotes the set of actions, ${\mathcal{R}}$ denotes the set of rewards, and $p:{\mathcal{S}}\times{\mathcal{R}}\times{\mathcal{S}}\times{\mathcal{A}}\to[0,1]$ denotes the transition dynamics. At time step $t$, the agent is in state $S_{t}\in{\mathcal{S}}$, takes action $A_{t}\in{\mathcal{A}}$ using a behavior policy $b:{\mathcal{A}}\times{\mathcal{S}}\to[0,1]$, observes the next state $S_{t+1}\in{\mathcal{S}}$ and reward $R_{t+1}\in{\mathcal{R}}$ according to the transition dynamics $p(s^{\prime},r\mid s,a)=\Pr(S_{t+1}=s^{\prime},R_{t}=r\mid S_{t}=s,A_{t}=a)$. We consider continuing problems, where the agent-environment interaction goes on ad infinitum. The agent’s goal is to maximize the average reward obtained over a long time (formally defined in (2)). We consider methods that try to achieve this goal by estimating the expected discounted sum of rewards from each state for $\gamma\in[0,1)$: $v_{\pi}^{\gamma}(s)\doteq\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}R_{t+1}\mid S_{t}=s,A_{t:\infty}\sim\pi],\forall s$. Here, the discount factor is not part of the problem but an algorithm parameter (see Naik et al. (2019) or Sutton & Barto’s (2018) Section 10.4 for an extended discussion on objectives for continuing problems).
Reward centering is a simple idea: subtract the empirical average of the observed rewards from the rewards. Doing so makes the rewards appear mean-centered. The effect of mean-centered rewards is well known in the bandit setting. For instance, Sutton and Barto (2018: Section 2.8) demonstrated that estimating and subtracting the average reward from the observed rewards can significantly improve the rate of learning. The benefits not only extend to the full RL problem, but are magnified as the discount factor $\gamma$ approaches one (bandit problems correspond to $\gamma=0$).
@@
and $e_{\pi}^{\gamma}(s)$ denotes an error term that go to zero as the discount factor goes to one (Blackwell, 1962: Theorem 4a; also see Puterman’s (1994) Corollary 8.2.4). This decomposition of the state values (1) also implies a similar decomposition for state–action values.
The constant state–action-independent term explains the improvements in the bandit setting (see Sutton & Barto’s (2018) Figure 2.5), where the action-value estimates are initialized to zero and the true values are centered around $+4$. The actions are selected based on their relative values, but each action-value estimate independently learns the constant offset. Approximation errors in estimating the offset can easily mask the relative differences in actions, especially if the offset is large.
In the full RL problem, the state(–action)-independent offset can indeed be quite large. For intuition, consider a three-state Markov reward process in Figure 2 induced by a policy $\pi$ in some MDP. There is a reward of $+3$ on going from state $s_{A}$ to $s_{B}$; 00 otherwise. The average reward $r(\pi)$ is $1$. The discounted state values for three discount factors are shown in Figure 2. Note the magnitude of the values and especially the jump when the discount factor is increased. Now consider the values with the constant offset subtracted from each state, $v_{\pi}^{\gamma}(s)-r(\pi)/(1-\gamma)$, which we call the centered discounted values: $\tilde{v}_{\pi}^{\gamma}$. The magnitudes of the centered values are much smaller, and differ only slightly when the discount factor is increased. The differential values are also shown for reference. These trends hold in general: for any problem, the magnitude of the discounted values shoots up as the discount factor approaches one (and may cause computational or numerical instability); meanwhile, the centered discounted values do not change much and approach the differential values.
$s_{A}$ $s_{B}$ $s_{C}$
@@
Figure 2: Comparison of the standard discounted values and the centered discounted values on a simple three-state problem.
@@
where $\gamma\in[0,1]$. When $\gamma=1$, the centered discounted values are the same as the differential values, that is, $\tilde{v}_{\pi}^{\gamma}(s)=\tilde{v}_{\pi}(s),\forall s$; when $r(\pi)=0,\tilde{v}_{\pi}^{\gamma}(s)=v_{\pi}^{\gamma}(s),\forall s$. More generally, the centered discounted values are the differential values plus the error terms from the Laurent-series decomposition (1).
Reward centering thus enables capturing all the information within the discounted value function via two components: (1) the constant average reward and (2) the centered discounted value function. Such a decomposition can be immensely valuable: (a) As $\gamma\to 1$, the discounted values tend to explode but the centered discounted values remain small and tractable. (b) If the problems’ rewards are shifted by a constant $c$, the magnitude of the discounted values increases by $c/(1-\gamma)$; the centered discounted values are unchanged because the average reward increases by $c$. Both of these effects are demonstrated in the following sections.
Reward centering also enables the design of algorithms in which the discount factor (an algorithm parameter) can be changed within the lifetime of a learning agent. This is usually inefficient or ineffective with standard algorithms since the uncentered discounted values can change massively (see, e.g., Figure 2). In contrast, the centered values may not change much, and the changes become minuscule as the discount factor approaches 1. We discuss this exciting direction in the final section.
To avail these potential benefits, we need to estimate the average reward from data. We shall now see that even the simplest method can take us quite far.
@@ 2 Simple Reward Centering
The simplest way to estimate the average reward is to maintain a running average of the rewards observed so far. That is, if $\bar{R}\in\mathbb{R}$ denotes the estimate of the average reward, after $t$ time steps, $\bar{R}_{t}=\sum_{k=1}^{t}R_{k}$. More generally, the estimate can be updated with a step-size parameter $\beta_{t}$:
@@
This update leads to an unbiased estimate of the average reward $r(\pi)$ corresponding to the policy $\pi$ with which the data is generated (that is, in the on-policy setting) if the step sizes follow the standard conditions (Robbins & Monro, 1951). To estimate the centered discounted values for a given policy $\pi$, we can use standard algorithms with the rewards centered using the current estimate of the average reward. For instance, after the transition $(S_{t},A_{t},R_{t+1},S_{t+1})$, the TD-learning algorithm can update its centered-value estimates $\tilde{V}^{\gamma}:{\mathcal{S}}\to\mathbb{R}$ using a step-size parameter $\alpha_{t}$ as:
@@
$\tilde{V}_{t+1}(S_{t})\doteq V_{t}(S_{t})+\alpha_{t}\big{[}(R_{t+1}-\bar{R}_{t})+\gamma\tilde{V}_{t}^{\gamma}(S_{t+1})-\tilde{V}_{t}^{\gamma}(S_{t})\big{]}.$ (5)
We call this on-policy algorithm that updates the value and average-reward estimates as in (4) and (5) as TD-learning with simple reward centering. As we demonstrate shortly, this simple approach to reward centering works quite well.
@@
Consider an MDP with seven states in a row with two actions in each state. The right action from the rightmost state leads to the middle state with reward $+7$ and the left action from the leftmost state leads to the middle state with reward $+1$; all other transitions have zero rewards. The target policy takes both actions in each state with equal probability, that is, $\pi(\text{left}|\cdot)=\pi(\text{right}|\cdot)=0.5$. The average reward corresponding to this policy is $0.25$. We tested three variants of the TD-learning algorithm: (a) standard TD-learning, (b) TD-learning with rewards that are centered by an oracle, (c) TD-learning with simple reward centering. We performed the same experiment with two discount factors, $\gamma=0.9$ and $0.99$, and evaluated the root mean-squared value error (RMSVE) of the estimates and the true discounted values at each time step w.r.t. the steady-state distribution of the states induced by the target policy $\pi$. The centering and oracle-centered methods estimate the centered discounted value function $\tilde{{\mathbf{v}}}^{\gamma}_{\pi}$, so for an apples-to-apples comparison, we added $\bar{R}_{t}/(1-\gamma)$ to the centered estimates to compute the uncentered value estimates at each time step.
The leftmost column of Figure 3 shows the learning curves for this on-policy experiment. The plots also contain the learning curves for another centering method that we shall discuss in the next section. In each plot, the curves correspond to the value of $\alpha$ that resulted in the fastest rate of learning for uncentered TD-learning baseline over the training period. For the centering methods, we chose a good value of $\eta$, where $\beta_{t}=\eta\alpha_{t}$ (without loss of generality), from a coarse search over a broad range. Each solid point represents the RMSVE averaged over the 50 independent runs; the shaded region denotes one standard error estimated from the data. Taken together, these plots help assess the average high-level trends that we are interested in. There is no separate training and testing period. All the experiments in this paper follow this general experimental setup; more details (including number of runs, range of parameter values tested) are in Appendix C.
@@
First, note that the learning curves start much lower when the rewards are centered by an oracle; for the other approaches, the first error is of the order $r(\pi)/(1-\gamma)$. Standard TD-learning (blue) eventually converges to the same error rate as the oracle-centered version (orange), as expected. Learning the average reward and subtracting it (green) indeed helps reduce the RMSVE much faster compared to when there is no centering. However, the eventual error rate is slightly higher, which is expected because the average-reward estimate is changing over time, leading to more variance in the updates compared to the uncentered or oracle-centered version. Similar trends hold for the larger discount factor (lower left), with the uncentered approach appearing much slower in comparison (note the difference in axes’ scales). In both cases, we verified that the average-reward estimate across the runs was around $0.25$.
@@
Limitations in the Off-policy Setting: (4) leads to an unbiased estimate of the behavior policy’s average reward, which means that in the off-policy setting, the average-reward estimate $\bar{R}$ will converge to $r(b)$, not to $r(\pi)$. Adding an importance-sampling ratio to the update is not enough to guarantee convergence to $r(\pi)$ because importance sampling only corrects the mismatch in action distributions, not the mismatch in the resulting state distributions.
@@
In summary, the effectiveness of reward centering increases with the accuracy of the average-reward estimate. Thus, even the simple method of reward centering (4) can be highly effective when the average reward of the behavior policy is close to that of the target policy. This may be true when the two policies are similar, like a greedy target policy and an $\epsilon$-greedy behavior policy with a relatively small value of $\epsilon$. However, the benefits of reward centering in terms of rate of learning may reduce and even disappear as the difference in the two policies increases. We shall now consider a subtly advanced approach to estimate the average reward more accurately in the off-policy setting.

### v2

@@ Abstract
We show that discounted methods for solving continuing reinforcement learning problems can perform significantly better if they center their rewards by subtracting out the rewards’ empirical average. The improvement is substantial at commonly used discount factors and increases further as the discount factor approaches one. In addition, we show that if a problem’s rewards are shifted by a constant, then standard methods perform much worse, whereas methods with reward centering are unaffected. Estimating the average reward is straightforward in the on-policy setting; we propose a slightly more sophisticated method for the off-policy setting. Reward centering is a general idea, so we expect almost every reinforcement-learning algorithm to benefit by the addition of reward centering.
@@
Figure 1: Learning curves showing the difference in performance of Q-learning with and without reward centering for different discount factors on the Access-Control Queuing problem (Sutton & Barto, 1998). Plotted is the average per-step reward obtained by the agent across 50 runs w.r.t. the number of time steps of interaction. The shaded region denotes one standard error. See Section 4.
In this paper, we show that the simple idea of estimating and subtracting the average reward from the observed rewards can lead to a significant improvement in performance (as in Figure 1) when using common discounted methods such as actor–critic methods (Barto et al., 1983) or Q-learning (Watkins & Dayan, 1992). The underlying theory dates back to 1962 with Blackwell’s seminal work on dynamic programming in discrete Markov decision processes (MDPs). We are still realizing some of its deeper implications, and we discuss the following two in particular:
@@
    Furthermore, mean-centering the rewards (unsurprisingly) makes standard methods robust to any constant offset in the rewards. This can be useful in reinforcement learning applications in which the reward signal is unknown or changing.
We begin with what reward centering is and why it can be beneficial (Section 1). We then show how reward centering can be done, starting with the simplest form (within the prediction problem), and show that it can be highly effective when used with discounted-reward temporal difference algorithms (Section 2). The off-policy setting requires more sophistication; for it we propose another way of reward centering based on recent advances in the average-reward formulation for reinforcement learning (Section 3). Next, we present a case study of using reward centering with Q-learning, in which we (a) propose a convergence result based on recent work by Devraj and Meyn (2021) and (b) showcase consistent trends across a series of control problems that require tabular, linear, and non-linear function approximation (Section 4). Finally, we discuss the limitations of the proposed methods and propose directions of future work (Section 5).
@@ 1 Theory of Reward Centering
We formalize the interaction between the agent and the environment by a finite MDP $({\mathcal{S}},{\mathcal{A}},{\mathcal{R}},p)$, where ${\mathcal{S}}$ denotes the set of states, ${\mathcal{A}}$ denotes the set of actions, ${\mathcal{R}}$ denotes the set of rewards, and $p:{\mathcal{S}}\times{\mathcal{R}}\times{\mathcal{S}}\times{\mathcal{A}}\to[0,1]$ denotes the transition dynamics. At time step $t$, the agent is in state $S_{t}\in{\mathcal{S}}$, takes action $A_{t}\in{\mathcal{A}}$ using a behavior policy $b:{\mathcal{A}}\times{\mathcal{S}}\to[0,1]$, observes the next state $S_{t+1}\in{\mathcal{S}}$ and reward $R_{t+1}\in{\mathcal{R}}$ according to the transition dynamics $p(s^{\prime},r\mid s,a)=\Pr(S_{t+1}=s^{\prime},R_{t+1}=r\mid S_{t}=s,A_{t}=a)$. We consider continuing problems, where the agent-environment interaction goes on ad infinitum. The agent’s goal is to maximize the average reward obtained over a long time (formally defined in (2)). We consider methods that try to achieve this goal by estimating the expected discounted sum of rewards from each state for $\gamma\in[0,1)$: $v_{\pi}^{\gamma}(s)\doteq\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}R_{t+1}\mid S_{t}=s,A_{t:\infty}\sim\pi],\forall s$. Here, the discount factor is not part of the problem but an algorithm parameter (see Naik et al. (2019) or Sutton & Barto’s (2018) Section 10.4 for an extended discussion on objectives for continuing problems).
Reward centering is a simple idea: subtract the empirical average of the observed rewards from the rewards. Doing so makes the modified rewards appear mean centered. The effect of mean-centered rewards is well known in the bandit setting. For instance, Sutton and Barto (2018, Section 2.8) demonstrated that estimating and subtracting the average reward from the observed rewards can significantly improve the rate of learning. Here we show that the benefits extend to the full reinforcement learning problem and are magnified as the discount factor $\gamma$ approaches one.
@@
and $e_{\pi}^{\gamma}(s)$ denotes an error term that goes to zero as the discount factor goes to one (Blackwell, 1962: Theorem 4a; also see Puterman’s (1994) Corollary 8.2.4). This decomposition of the state values also implies a similar decomposition for state–action values.
The Laurent-series decomposition explains how reward centering can help learning in bandit problems such as the one in Sutton & Barto’s (2018) Figure 2.5. There, the action-value estimates are initialized to zero and the true values are centered around $+4$. The actions are selected based on their relative values, but each action-value estimate must independently learn the same constant offset. Approximation errors in estimating the offset can easily mask the relative differences in actions, especially if the offset is large.
In the full reinforcement learning problem, the state-independent offset can be quite large. For example, consider the three-state Markov reward process shown Figure 2 (induced by some policy $\pi$ in some MDP). The reward is $+3$ on transition from state A to state B, and 00 otherwise. The average reward is $r(\pi)=1$. The discounted state values for three discount factors are shown in the table. Note the magnitude of the standard discounted values and especially the jump when the discount factor is increased. Now consider the discounted values with the constant offset subtracted from each state, $v_{\pi}^{\gamma}(s)-r(\pi)/(1-\gamma)$, which we call the centered discounted values. The centered values are much smaller in magnitude and change only slightly when the discount factor is increased. The differential values are also shown for reference. These trends hold in general: for any problem, the magnitude of the discounted values increase dramatically as the discount factor approaches one whereas the centered discounted values change little and approach the differential values.
State A B C
@@
Figure 2: Comparison of the standard and the centered discounted values on a simple example.
@@
where $\gamma\in[0,1]$. When $\gamma=1$, the centered discounted values are the same as the differential values, that is, $\tilde{v}_{\pi}^{\gamma}(s)=\tilde{v}_{\pi}(s),\forall s$. More generally, the centered discounted values are the differential values plus the error terms from the Laurent-series decomposition, as shown on the right above.
Reward centering thus enables capturing all the information within the discounted value function via two components: (1) the constant average reward and (2) the centered discounted value function. Such a decomposition can be immensely valuable: (a) As $\gamma\to 1$, the discounted values tend to explode but the centered discounted values remain small and tractable. (b) If the problems’ rewards are shifted by a constant $c$, then the magnitude of the discounted values increases by $c/(1-\gamma)$, but the centered discounted values are unchanged because the average reward increases by $c$. These effects are demonstrated in the following sections.
Reward centering also enables the design of algorithms in which the discount factor (an algorithm parameter) can be changed within the lifetime of a learning agent. This is usually inefficient or ineffective with standard discounted algorithms because their uncentered values can change massively (Figure 2). In contrast, centered values may change little, and the changes become minuscule as the discount factor approaches 1. We discuss this exciting direction in the final section.
To obtain these potential benefits, we need to estimate the average reward from data. In the next section we show that even the simplest method can be quite effective.
@@ 2 Simple Reward Centering
The simplest way to estimate the average reward is to maintain a running average of the rewards observed so far. That is, if $\bar{R}_{t}\in\mathbb{R}$ denotes the estimate of the average reward after $t$ time steps, then $\bar{R}_{t}=\sum_{k=1}^{t}R_{k}$. More generally, the estimate can be updated with a step-size parameter $\beta_{t}$:
@@
This update leads to an unbiased estimate of the average reward $\bar{R}_{t}\approx r(\pi)$, for the policy $\pi$ generating the data, if the step sizes follow standard conditions (Robbins & Monro, 1951).
Simple centering (4) can be used with almost any reinforcement learning algorithm. For example, it can be combined with conventional temporal-difference (TD) learning (see Sutton, 1988a) to learn a state-value function estimate $\tilde{V}^{\gamma}:{\mathcal{S}}\to\mathbb{R}$ by updating, on transition from $t$ to $t+1$:
$\tilde{V}^{\gamma}_{t+1}(S_{t})\doteq\tilde{V}^{\gamma}_{t}(S_{t})+\alpha_{t}\big{[}(R_{t+1}-\bar{R}_{t})+\gamma\tilde{V}_{t}^{\gamma}(S_{t+1})-\tilde{V}_{t}^{\gamma}(S_{t})\big{]},$ (5)
@@
with $\tilde{V}^{\gamma}_{t+1}(s)\doteq\tilde{V}^{\gamma}_{t}(s),\forall s\neq S_{t}$, where $\alpha_{t}>0$ is a step-size parameter.
We used four algorithmic variations of (5) differing only in the definition of $\bar{R}_{t}$ in our first set of experiments. One algorithm used $\bar{R}_{t}=0,\forall t$, and thus involves no reward centering. The second algorithm used the best possible estimate of the average reward: $\bar{R}_{t}=r(\pi),\forall t$; we call this oracle centering. The third algorithm used simple reward centering as in (4). The fourth algorithm used a more sophisticated kind of reward centering which we discuss in the next section.
@@
The environment was an MDP with seven states in a row with two actions in each state. The right action from the rightmost state leads to the middle state with a reward of $+7$ and the left action from the leftmost state leads to the middle state with a reward of $+1$; all other transitions have zero rewards. The target policy takes both actions in each state with equal probability, that is, $\pi(\text{left}|\cdot)=\pi(\text{right}|\cdot)=0.5$. The average reward corresponding to this policy is $r(\pi)=0.25$.
Our first experiment applied the four algorithms to the seven-state MDP with two discount factors, $\gamma=0.9$ and $0.99$. All algorithms were run with a range of values for the step-size parameters $\alpha$. The algorithms that learned to center were run with different values of $\eta$, where $\beta=\eta\alpha$ (without loss of generality). Each parameter setting for each algorithm was run for 50,000 time steps, and then repeated for 50 runs. The full experimental details are in Appendix C. As a measure of performance at time $t$, we used the root-mean-squared value error (RMSVE; see Sutton & Barto, 2018, Section 9.2) between $\tilde{V}^{\gamma}_{t}$ and $\tilde{v}_{\pi}^{\gamma}$ for the centered algorithms, and between $\tilde{V}^{\gamma}_{t}$ and $v_{\pi}^{\gamma}$ for the algorithm without centering. There was no separate training and testing period.
Learning curves for this experiment and each value of $\gamma$ are shown in the first column of Figure 3. For all algorithms, we show only curves for the $\alpha$ value that was best for TD-learning without reward centering. For the centering methods, the curve shown is for the best choice of $\eta$ from a coarse search over a broad range. Each solid point represents the RMSVE averaged over the 50 independent runs; the shaded region shows one standard error.
@@
First note that the learning curves start much lower when the rewards are centered by an oracle; for the other algorithms, the first error is of the order $r(\pi)/(1-\gamma)$. TD-learning without centering (blue) eventually reached the same error rate as the oracle-centered algorithm (orange), as expected. Learning the average reward and subtracting it (green) indeed helps reduce the RMSVE much faster compared to when there is no centering. However, the eventual error rate is slightly higher, which is expected because the average-reward estimate is changing over time, leading to more variance in the updates compared to the uncentered or oracle-centered version. Similar trends hold for the larger discount factor (lower left), with the uncentered approach appearing much slower in comparison (note the difference in axes’ scales). In both cases, we verified that the average-reward estimate across the runs was around $0.25$.
@@
Limitations in the Off-policy Setting: (4) leads to an unbiased estimate of the behavior policy’s average reward, which means that in the off-policy setting the average-reward estimate $\bar{R}$ will converge to $r(b)$, not $r(\pi)$. Adding an importance-sampling ratio to the update is not enough to guarantee convergence to $r(\pi)$ because importance sampling only corrects the mismatch in action distributions, not the mismatch in the resulting state distributions.
@@
In summary, the effectiveness of reward centering increases with the accuracy of the average-reward estimate. Thus, even the simple method of reward centering (4) can be highly effective when the average reward of the behavior policy is close to that of the target policy. This may be true when the two policies are similar, like a greedy target policy and an $\epsilon$-greedy behavior policy with a relatively small value of $\epsilon$. However, the benefits of reward centering in terms of rate of learning may reduce and even disappear as the difference in the two policies increases. In the following section, we present a subtly advanced approach to estimate the average reward more accurately in the off-policy setting.